A man travels for work averaging 40mph. On the return trip he averages 50mph, and it takes him an hour less time. How long was the return trip?
Answer by rwm(914)

Same as the original trip.
He returned.
Is this a trick question?
r*t=d
40*t=50(t-1)
40t=50t-50
50=10t
5=t
check
40*5=50*4
200=200
ok
200 miles either way